OpenStudy (anonymous):

Find the consecutive even number such that the sum of 3/5 of first, 1/2 of second and 3/8 of the third is 32 www.teachon.tk

OpenStudy (lgbasallote):

x = first number x+2 = second number x+4 = 3rd number 3/5 (x) + 1/2 (x+2) + 3/8 (x+4) = 32 i assume you can take over from here